OpenStudy (anonymous):

Fanning + medal What is the Greatest Common Factor of x4 and x3? x x2 x3 x4

OpenStudy (anonymous):

I believe you were already given this answer. Do you have a question about it?

OpenStudy (anonymous):

Someone says its x^3 is it that answer ? @gypsy1274

OpenStudy (anonymous):

It is. Do you know why?

OpenStudy (anonymous):

Oh i thought it wouldn't its okay thanks gypsy

OpenStudy (goldphenoix):

It is the lowest exponent out of the 2 terms. Why is that? Well, pretend x is 2. 2^3 = 2 *2 *2 2^4 = 2 *2 *2 *2 You want to look for the greatest common factor. So is 2 a factor of those 2? Yes. But is the greatest? Well, what about 2^2? 2^2 = 4. Is 4 the greatest common factor? What about, 2^3? 2^3 = 2 *2 *2 = 8. Is that the greatest common factor? What about 2^4. 2^4 = 2 * 2 *2 *2 = 16. 16 is greater than 2^3, or 8. Therefore, that cannot be a factor of it. :)

OpenStudy (anonymous):

goldphenoix im about to close the question i dont want to be rude so ill wait

OpenStudy (goldphenoix):

I finished 10 seconds ago. :)

OpenStudy (goldphenoix):

Did you understand my explanation?

OpenStudy (anonymous):

Theres no explanation

OpenStudy (goldphenoix):

It is the lowest exponent out of the 2 terms. Why is that? Well, pretend x is 2. 2^3 = 2 *2 *2 2^4 = 2 *2 *2 *2 You want to look for the greatest common factor. So is 2 a factor of those 2? Yes. But is the greatest? Well, what about 2^2? 2^2 = 4. Is 4 the greatest common factor? What about, 2^3? 2^3 = 2 *2 *2 = 8. Is that the greatest common factor? What about 2^4. 2^4 = 2 * 2 *2 *2 = 16. 16 is greater than 2^3, or 8. Therefore, that cannot be a factor of it. :)

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

x^3 right?

OpenStudy (goldphenoix):

Sorry about it being unorganized. It was organized in my first reply. Yes it is, x^3 since it is the lowest exponent out of the 2 terms. :)

OpenStudy (anonymous):

Lol your fine im going to be putting up more questions look for mines :)